OpenStudy (anonymous):

Find dy/dx of y=(x²-7)/(x^1/3). It says I shouldn't have to use the quotient rule and that the final answer should just be in the form ax^m+bx^n. Could someone help me figure this out? Thanks.

OpenStudy (anonymous):

Sure, just divide each term in the numerator by the denominator. \[ \frac{(a+b) }{c} = \frac{a}{c} + \frac{b}{c} \]

OpenStudy (anonymous):

That makes more sense. How do you find dy/dx for the exponent that is in the denominator though?

OpenStudy (anonymous):

Once fractions get involved, I kinda screw up on these things

OpenStudy (anonymous):

Fractions and exponents in the denominator, haha

OpenStudy (anonymous):

So that would simplify to \[ \frac{x^2-7}{x^{1/3}} = \frac{x^2}{x^{1/3}} - \frac{7}{x^{1/3}} = x^{2/3} - 7x^{-1/3}\]

OpenStudy (anonymous):

why would it be x to the 2/3 power?

OpenStudy (anonymous):

Oops, I'm sorry, that should be 5/3, not 2/3

OpenStudy (anonymous):

Ah ok that makes sense now. Thanks :)

OpenStudy (anonymous):

And then I just need to find dy/dx of that, right?

OpenStudy (anonymous):

Right.

OpenStudy (anonymous):

Thank you!